OpenStudy (anonymous):

HELP PLEASE!! Determine the exact value of x. a.32x = 5(3x) + 36 b.(1/8)x - 3 = 2 × 162x + 1 c.3x² + 20 = (1/27)3x

OpenStudy (phi):

b. is a good example of why you should not use x to mean multiplication. Use a dot instead.

OpenStudy (anonymous):

A.) Mulitply 5 by 3 to get 15: 32x=15x + 36 Move all variable terms from the right side of the equal sign to the left. Expand 15x. 32x-15x=15x + 36 - 15x. Combine Like Terms. 17x=36 Divide both sides by 17. Dividing by 17 undoes the multiplication by 17. x=36/17